BANDALA HAS 30 OUNCES OF 60% SULFURIC ACID SOLUTION. HOW MANY OUNCES OF 15% SULFURIC ACID SOLUTION SHOULD HE ADD TO OBTAIN A SOLUTION THAT IS 10% SULFURIC ACID SOLUTION?
PLEASE HELP ME.

Answer by mananth(16946)

BANDALA HAS 30 OUNCES OF 60% SULFURIC ACID SOLUTION. HOW MANY OUNCES OF 15% SULFURIC ACID SOLUTION SHOULD HE ADD TO OBTAIN A SOLUTION THAT IS 10% SULFURIC ACID SOLUTION?
This is not possible. the concentration of the diluting solution is higher than the mixture concentration. Check the problem or give this explanation.
..
60% SULFURIC ACID SOLUTION --------------30
15% SULFURIC ACID SOLUTION ---------------x
10% SULFURIC ACID SOLUTION?--------------(x+30)
60*30 +15*x=10(x+30)
1800+15x=10x+300
10x-15x= 1800-300
-5x=1500
/-5
-300 This is absurd. why ?
..
You need a solution which is lower than the concentrations of the solutions mixed. Hence this is not possible.